OpenStudy (anonymous):

Answer all ?'s and receive a medal :)! v=-0.0098t+c, c is exhaust velocity in km/s, in R where c is the exhaust velocity in km/s, t is firing time, R is mass ratio. Needs to reach 7.7 km/s in order to orbit at 300 km. 1. If rocket has mass ratio of 18, exhaust velocity of 2.2 km/s and firing time is 25s. Does it make orbit? 2. What is the mass velocity in order to orbit if it has exhaust velocity of 2.5 km/s and firing time of 29s 3. Exhaust velocity- 2.4 km/s, firing time 28s, max velocity of 7.8 km/s. What is the mass ratio?

OpenStudy (whpalmer4):

it looks to me like you didn't copy the problem exactly...

OpenStudy (whpalmer4):

" km/s, in R where c i" is my tipoff

OpenStudy (whpalmer4):

"v=-0.0098t+c, c is exhaust velocity in km/s, in R where c is the exhaust velocity " does not make sense.

OpenStudy (anonymous):

so you cant solve it?

OpenStudy (whpalmer4):

key to solving a problem is knowing what the problem is. you've given a scrambled formula. the correct formula is \[v = -0.0098t+c \ln R\]

OpenStudy (whpalmer4):

problem 2 is the mass ratio, not mass velocity

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, I've solved all 3. How about you? :-)

OpenStudy (whpalmer4):

"ln" is the natural logarithm, or logarithm to the base e = 2.718281828...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

did you get 6.114 for 1?

OpenStudy (whpalmer4):

6.11382, yep.

OpenStudy (whpalmer4):

so that's a no-go on orbit.

OpenStudy (anonymous):

im at .1225=2.5lnR what now?

OpenStudy (whpalmer4):

how did you get that? that's going to give you R = 1.05 or so, which is not reasonable

OpenStudy (anonymous):

can you show me how to do it?

OpenStudy (whpalmer4):

2. What is the mass velocity in order to orbit if it has exhaust velocity of 2.5 km/s and firing time of 29s \[v = -0.0098t + c \ln R\]\[7.7 = -0.0098(29) + 2.5\ln R\]\[7.7 = -0.2842+2.5\ln R\]\[7.4158 = 2.5\ln R\]\[2.96632 = \ln R\]\[e^{2.96632} = R\]

OpenStudy (whpalmer4):

hang on, let me check that...

OpenStudy (whpalmer4):

nuts, I subtracted when I should have added! \[7.7 = -0.2842+2.5\ln R\]\[7.9842=2.5\ln R\]\[3.19368=\ln R\]\[R=e^{3.19368}\]

OpenStudy (anonymous):

how about 3

OpenStudy (whpalmer4):

do you have an answer to check?

OpenStudy (anonymous):

27.73

OpenStudy (whpalmer4):

Hmm. 3. Exhaust velocity- 2.4 km/s, firing time 28s, max velocity of 7.8 km/s. What is the mass ratio? \[7.8 = -0.0098(28)+2.4\ln R\]\[8.0744 = 2.4\ln R\]\[3.36433=\ln R\]

OpenStudy (whpalmer4):

that's a different result than you got.

OpenStudy (whpalmer4):

Ah, you used 7.7 instead of 7.8

OpenStudy (anonymous):

should i have used 7.8?

OpenStudy (whpalmer4):

Well, the problem does say "max velocity of 7.8 km/s"